OpenStudy (anonymous):

Given the point (3, 4) and the slope of 6, find y when x = 27. 122 14 141 148

OpenStudy (anonymous):

?

OpenStudy (anonymous):

141

OpenStudy (lucaz):

you have to use the point-slope form

OpenStudy (lucaz):

y-y1=m(x-x1), you the slope m, the x1 and y1, plug them in and you find y

OpenStudy (lucaz):

also x, wich is 27

OpenStudy (lucaz):

(y-(4))=(6)(27-(3)) y-4=144 y=148